# Improving Dialogue Agents by Decomposing One Global Explicit Annotation with Local Implicit Multimodal Feedback

### v1

@@ Abstract
We describe an approach for aligning an LLM-based dialogue agent based on global (i.e., dialogue-level) rewards, while also taking into account naturally-occurring multimodal signals. At a high level, our approach (dubbed GELI [Uncaptioned image] ) learns a local, turn-level reward model by decomposing the human-provided Global Explicit (GE) session-level reward, using Local Implicit (LI) multimodal reward signals to crossmodally shape the reward decomposition step. This decomposed reward model is then used as part of the standard RHLF pipeline improve an LLM-based dialog agent. We run quantitative and qualitative human studies to evaluate the performance of our GELI approach, and find that it shows consistent improvements across various conversational metrics compared to baseline methods.
@@
[…]
@@ 1 Introduction
Developing social dialogue agents that can interact and collaborate with humans over a long horizon remains a longstanding goal of artificial intelligence. Large language models (LLM) pretrained at scale on the next-word prediction objective and then subsequently aligned to human preference via RLHF represent a significant step in this direction Ouyang et al. (2022), even leading to successful commercial applications.
@@
In this paper, we describe a joint framework called GELI [Uncaptioned image] , which integrates global explicit (GE) and local implicit (LI) feedback. GELI makes it possible to align an LLM-based dialogue agent based on global rewards, while simultaneously taking into account naturally-occurring multimodal signals. Our formulation brings together the idea of training a reward model which decomposes a single global explicit annotation score that is shaped by local implicit multimodal signals, which is subsequently used to align an LLM-based dialogue agent via RLHF. Specifically, we use GELI to learn a reward function based on the overall affect of the user (i.e., how positive the user felt at the end of the conversation) from a large-scale long-horizon multimodal dialogue dataset Reece et al. (2023). Our local implicit multimodal signal comes from an affect classifer based on facial expression. We find that the reward function learned via GELI can be used train a dialogue agent that has improved ability across various metrics of conversational quality including sensibleness, reusability, and specificity Lee et al. (2022).
@@ 2 Related Works
Figure 1: Overview of our proposed method: GELI Refer to caption . Left: The reward function training involves decomposing a single global explicit (GE) feedback, with the guidance of multimodal local implicit (LI) feedback, such as visual facial affect. Right: We utilize the decomposed reward function to update the language model, where the language model generates utterances and the reward function assigns a score to be optimized via PPO.
@@
Incorporating human preference feedback into a reward model, and subsequently optimizing a language model to output text that reward model scores highly with an RL algorithm, has been shown to result in language models that generate outputs humans generally prefer (Ouyang et al., 2022). This process has been applied to summarization (Ziegler et al., 2019; Stiennon et al., 2020; Wu et al., 2021), answering questions with long-form answers using text retrieved from the web (Nakano et al., 2021; Menick et al., 2022), generating engaging responses in a dialogue settings (Thoppilan et al., 2022; Cohen et al., 2022) and following human instructions (Kojima et al., 2021; Suhr and Artzi, 2022; Kim et al., 2023). However, these methods generally require collecting fine-grained annotations for each generated response to train the reward function, which is difficult to obtain at scale for long-horizon dialogue.
@@ 3 Background
4 Methods: GELI [Uncaptioned image]
The reward function $r_{\theta}$ in standard adaptation techniques relies on intermediate fine-grained annotations, requiring manual human annotations at each generated utterance. However, in many long-term dialogue settings there is only a single global explicit (GE) annotated reward for each session. Given a trajectory of the multi-turn dialogue $\tau$, the global explicit reward $R_{GE}(\tau)$ is a scalar reward at the end of the interaction, such as how positively the user felt about the conversation. This GE reward can be decomposed via sum decomposition (more details in Sec. 4.1) with the GE loss function $\mathcal{L}_{\text{GE}}$. A core novelty of our proposed GELI [Uncaptioned image] approach is that the decomposition of the GE reward will be guided by some Local Implicit (LI) feedback. Concretely, in many dialog applications/datasets of interest there are rich multimodal signals, which is can provide intermediate signals that are useful for the decomposition of the single global explicit reward. We thus perform cross-modal distillation of the signals from such multimodal signals into the individually decomposed text-only reward function via the LI loss function $\mathcal{L}_{\text{LI}}$ (more details in Sec. 4.2).
Figure 2: Example of GELI Refer to caption reward score predictions for an unseen conversation from the dataset. Top left: Reward scores unrolled over an unseen conversation, where the mean is subtracted. We examine a random sampled snippet, where we find that our decomposed reward function assigns higher values to meaningful utterances.
@@ 4.1 GE: Decomposing One Global Explicit Annotation
Global explicit reward is a human annotation at the end of the interaction, which can come in the form of a post-interaction score. Let $\tau$ denote the trajectory of the episode, i.e. $\tau=\left\langle s_{0},a_{0},s_{1},a_{1}\cdots,s_{T},a_{T}\right\rangle$. This reward represents the overall reward of trajectory $\tau$, $R_{\mathrm{GE}}(\tau)$. The agent in this episodic reinforcement learning paradigm must maximize the expected global explicit reward at the end of the conversation. One way to approximate the global explicit reward $R_{\mathrm{ep}}(\tau)$ is by sum decomposition via considering the sum of $r_{\theta}(s_{t},a_{t})$, across all the previous states $s_{t}$ and newly generated $a_{t}$:
@@
Then, this idea of sum-based return decomposition (RD), can be implemented via a least-squares-based approach, where the reward distribution is given by a learnt reward function, decomposing the episodic reward $R_{\mathrm{ep}}(\tau)$ in an additive way Arjona-Medina et al. (2019).
@@ 5 Experiments
In this section, we describe our experiments to evaluate our proposed GELI [Uncaptioned image] framework which performs reward function training with global explicit reward decomposition and local implicit visual feedback. All experiments are performed by (1) first, training a reward function (e.g. using GELI or one of its ablation variant only GE or only LI) (2) and use the trained reward functions in a reinforcement learning setup with PPO Schulman et al. (2017) to adapt the language model in generating better conversational responses. Due to computational resources, the training of reward functions and adaptations are performed over a single run.
@@ 6.1 Reward Function
Baselines Connection Positivity Social Inclination Interestingness Reuse Specific Sensible GELI Score
@@
LLAMA2 + GELI: RRD+VA 39.67 $\pm$ 7.32 44.33 $\pm$ 12.23 35.33 $\pm$ 10.87 37.33 $\pm$ 6.85 38.0 $\pm$ 10.2 41.67 $\pm$ 7.04 80.33 $\pm$ 4.5 80.67 $\pm$ 10.5 0.5419
Table 2: Human evaluation results on 100 samples for 3 seeds for 8 preference metrics where mean and std. are reported. Green indicates best score, Bold indicates scores that fall in the confidence interval of the best performing score. GELI [Uncaptioned image] performs better on 6 out of 8 metrics (emotional connection, positivity, social understanding, inclination, interestingness, reuse) and comparably to the best performing model on the other 2 metrics: specific and sensible.
@@
We refer the readers to the bottom of Table 1. The results are shown for the reward decomposition and visual feedback for the reward function trained with GELI: global explicit reward decomposition informed by local implicit multimodal feedback shaping. We find that the combination of random return decomposition (RRD) and visual affect (VA) achieves the best of both worlds, where the reward decomposition loss is low ($L_{GE}\downarrow$) and there is a noticeable difference in the reward function scores between positive and non-positive facial affect samples ($\Delta\hat{r}_{LI}>0$).
@@
6.2 Human Evaluation on Adapted LLM
We refer the reader to Table 2, where we find that the LLAMA-2 model with GELI outperforms all other approaches in most evaluation metrics and performs comparably with other baselines otherwise. Compared to base LLAMA-2, we see that there is a significant improvement in the level of emotional connection (+9%), positivity (+18%), understanding of social context (+10%), how interesting the responses are (+14%). It is especially impressive to note that there is a statisical difference in how inclined people wanted to talk to our model over others (+11%), and how much they would want to reuse our chatbot again (+14%). Interestingly, we see statistically signficant results for positivity, which is the most closely related to our primary optimization objective overall-affect, and inclination, reuse, which indicates which chatbot the User would speak to again.
@@ 7 Conclusion
We introduce GELI [Uncaptioned image] , which automatically decomposes a single Global Explicit post-interaction score, incorporating Local Implicit feedback from multimodal behaviors. The reward function trained via GELI is designed to align and improve the conversational capabilities of a language model. GELI performs global alignment of multi-turned interactions by locally rewarding parts of the interaction, shaped by multimodal local implicit feedback. Our proposed approach complements previous alignment approaches, such as RLHF, which requires fine-grained manual reward annotations. We run quantitative and qualitative human studies to evaluate the performance of our GELI approach, with results showing consistent performance boosts across conversational metrics.

### v2

@@ Abstract
We describe an approach for aligning an LLM-based dialogue agent based on global (i.e., dialogue-level) rewards, while also taking into account naturally-occurring multimodal signals. At a high level, our approach (dubbed GELI) learns a local, turn-level reward model by decomposing the human-provided Global Explicit (GE) session-level reward, using Local Implicit (LI) multimodal reward signals to crossmodally shape the reward decomposition step. This decomposed reward model is then used as part of the standard RLHF pipeline improve an LLM-based dialog agent. We run quantitative and qualitative human studies to evaluate the performance of our GELI approach, and find that it shows consistent improvements across various conversational metrics compared to baseline methods.
@@
Dong Won Lee1  Hae Won Park1  Yoon Kim1  Cynthia Breazeal1  Louis-Philippe Morency2 Massachusetts Institute of Technology1, Carnegie Mellon University2 dongwonl@mit.edu
@@ 1 Introduction
Developing social dialogue agents that can interact and collaborate with humans over a long horizon remains a longstanding goal of artificial intelligence. Large language models (LLM) pretrained at scale on the next-word prediction objective and then subsequently aligned to human preference via RLHF (Reinforcement with Human Feedback) represent a significant step in this direction Ouyang et al. (2022), even leading to successful commercial applications.
@@
In this paper, we describe a joint framework called GELI, which integrates global explicit (GE) and local implicit (LI) feedback. GELI makes it possible to align an LLM-based dialogue agent based on global rewards, while simultaneously taking into account naturally-occurring multimodal signals. Our formulation brings together the idea of training a reward model which decomposes a single global explicit annotation score that is shaped by local implicit multimodal signals, which is subsequently used to align an LLM-based dialogue agent via RLHF. Specifically, we use GELI to learn a reward function based on the overall affect of the user (i.e., how positive the user felt at the end of the conversation) from a large-scale long-horizon multimodal dialogue dataset Reece et al. (2023). Our local implicit multimodal signal comes from an affect classifer based on facial expression. We find that the reward function learned via GELI can be used train a dialogue agent that has improved ability across various metrics of conversational quality including sensibleness, reusability, and specificity Lee et al. (2022).
@@ 2 Related Works
Figure 1: Overview of our proposed method: GELI. Left: The reward function training involves decomposing a single global explicit (GE) feedback, with the guidance of multimodal local implicit (LI) feedback, such as visual facial affect. Right: We utilize the decomposed reward function to update the language model, where the language model generates utterances and the reward function assigns a score to be optimized via PPO Schulman et al. (2017).
@@
Incorporating human preference feedback into a reward model, and subsequently optimizing a language model to output text that reward model scores highly with an RL algorithm, has been shown to result in language models that generate outputs humans generally prefer (Ouyang et al., 2022). This process has been applied to summarization (Ziegler et al., 2019; Stiennon et al., 2020; Wu et al., 2021), answering questions with long-form answers using text retrieved from the web (Nakano et al., 2021; Menick et al., 2022), generating engaging responses in a dialogue settings (Thoppilan et al., 2022; Cohen et al., 2022) and following human instructions (Kojima et al., 2021; Suhr and Artzi, 2022; Kim et al., 2023b). However, these methods generally require collecting fine-grained annotations for each generated response to train the reward function, which is difficult to obtain at scale for long-horizon dialogue.
Utilizing Implicit Signals for Dialogue Agents
Many previous work utilize local implicit signals found only in the text, such as existence of next human turn, next human turn length, mean conversation length, sentiment and reaction in the next human utterance, retry rate, retention rate, or user rating Pang et al. (2023); Irvine et al. (2023). In contrary, ours is the first (1) to additionally utilize multimodal signals, and (2) use global signals in conjunction with the local implicit signals, which has been a crucial finding that contributed significantly to the performance boost in the human evaluation.
@@ 3 Background
4 Methods: GELI
The reward function $r_{\theta}$ in standard adaptation techniques relies on intermediate fine-grained annotations, requiring manual human annotations at each generated utterance. However, in many long-term dialogue settings there is only a single global explicit (GE) annotated reward for each session. Given a trajectory of the multi-turn dialogue $\tau$, the global explicit reward $R_{GE}(\tau)$ is a scalar reward at the end of the interaction, such as how positively the user felt about the conversation. This GE reward can be decomposed via sum decomposition (more details in Sec. 4.1) with the GE loss function $\mathcal{L}_{\text{GE}}$. A core novelty of our proposed GELI approach is that the decomposition of the GE reward will be guided by some Local Implicit (LI) feedback. Concretely, in many dialog applications/datasets of interest there are rich multimodal signals, which is can provide intermediate signals that are useful for the decomposition of the single global explicit reward. We thus perform cross-modal distillation of the signals from such multimodal signals into the individually decomposed text-only reward function via the LI loss function $\mathcal{L}_{\text{LI}}$ (more details in Sec. 4.2).
Figure 2: Example of GELI reward score predictions for an unseen conversation from the dataset. Top left: Reward scores unrolled over an unseen conversation, where the mean is subtracted. We examine a random sampled snippet, where we find that our decomposed reward function assigns higher values to meaningful utterances.
@@ 4.1 GE: Decomposing One Global Explicit Annotation
Global explicit reward is a human annotation at the end of the interaction, which can come in the form of a post-interaction score. Let $\tau$ denote the trajectory of the episode, i.e. $\tau=\left\langle s_{0},a_{0},s_{1},a_{1}\cdots,s_{T},a_{T}\right\rangle$. This reward represents the overall reward of trajectory $\tau$, $R_{\mathrm{GE}}(\tau)$. The agent in this episodic reinforcement learning paradigm must maximize the expected global explicit reward at the end of the conversation. One way to approximate the global explicit reward $R_{\mathrm{GE}}(\tau)$ is by sum decomposition via considering the sum of $r_{\theta}(s_{t},a_{t})$, across all the previous states $s_{t}$ and newly generated $a_{t}$:
@@
Then, this idea of sum-based return decomposition (RD), can be implemented via a least-squares-based approach, where the reward distribution is given by a learnt reward function, decomposing the episodic reward $R_{\mathrm{GE}}(\tau)$ in an additive way Arjona-Medina et al. (2019).
@@ 5 Experiments
In this section, we describe our experiments to evaluate our proposed GELI framework which performs reward function training with global explicit reward decomposition and local implicit visual feedback. All experiments are performed by (1) first, training a reward function (e.g. using GELI or one of its ablation variant only GE or only LI) (2) and use the trained reward functions in a reinforcement learning setup with PPO Schulman et al. (2017) to adapt the language model in generating better conversational responses. Due to computational resources, the training of reward functions and adaptations are performed over a single run.
@@ 6.1 Reward Function
CANDOR Reece et al. (2023) Connection Positivity Social Inclination Interestingness Reuse Specific Sensible GELI Score
@@
LLAMA2 + GELI: RRD+VA (Ours) 39.67 $\pm$ 7.32 44.33 $\pm$ 12.23 35.33 $\pm$ 10.87 37.33 $\pm$ 6.85 38.0 $\pm$ 10.2 41.67 $\pm$ 7.04 80.33 $\pm$ 4.5 80.67 $\pm$ 10.5 0.5419
Table 2: Human evaluation results on 100 samples for 3 seeds for 8 preference metrics where mean and std. are reported. Green indicates best score. GELI [Uncaptioned image] performs better on 6 out of 8 metrics (emotional connection, positivity, social understanding, inclination, interestingness, reuse) and comparably to the best performing model on the other 2 metrics: specific and sensible.
SODA Kim et al. (2023a) Connection Positivity Social Inclination Interestingness Reuse Specific Sensible
(/100%) ↑
GPT-3.5 (text-davinci-002) 40.1 ± 7.56 43.05 ± 3.4 48.13 ± 9.08 46.05 ± 3.44 49.11 ± 7.69 44.03 ± 2.01 78.14 ± 9.49 80.07 ± 7.72
LLAMA2 66.04 ± 4.79 70.0 ± 2.51 71.99 ± 6.28 67.0 ± 0.46 55.05 ± 8.24 65.99 ± 6.3 89.04 ± 2.65 89.99 ± 3.81
LLAMA2 + GE: RRD 30.98 ± 2.66 30.98 ± 5.04 34.04 ± 3.28 27.0 ± 7.43 24.98 ± 2.69 30.0 ± 2.51 43.97 ± 3.3 47.06 ± 4.34
LLAMA2 + LI: LS 62.0 ± 3.71 70.06 ± 4.52 75.02 ± 5.06 68.04 ± 3.41 59.0 ± 1.24 68.01 ± 3.72 86.04 ± 2.61 92.99 ± 1.47
LLAMA2 + LI: VA 55.02 ± 1.92 57.1 ± 7.21 63.04 ± 4.76 51.99 ± 0.67 43.97 ± 3.3 51.04 ± 3.08 76.03 ± 2.16 82.0 ± 2.49
LLAMA2 + GELI: RRD + VA (Ours) 71.01 ± 1.27 73.98 ± 1.76 76.98 ± 3.01 71.99 ± 1.65 66.97 ± 6.69 70.0 ± 2.51 90.02 ± 7.53 88.06 ± 4.73
Table 3: Human evaluation results on an unseen dataset, SODA Kim et al. (2023a), during all learning steps to demonstrate generalizability across datasets and dialogue scenarios. 33 samples for 3 seeds for 8 preference metrics where mean and std. are reported. Green indicates best score, 7 out of 8 metrics (emotional connection, positivity, social understanding, inclination, interestingness, reuse) and comparably to the best performing model on the other 1 metrics: specific and sensible.
@@
We refer the readers to the bottom of Table 1. The results are shown for the reward decomposition and visual feedback for the reward function trained with GELI: global explicit reward decomposition informed by local implicit multimodal feedback shaping. We find that the combination of random return decomposition (RRD) and visual affect (VA) achieves the best of both worlds.
It is important to look at both error metrics (GE and LI): the $L_{GE}$ metric is evaluating performance globally, comparing the final predicted score of the whole conversation with the ground truth (which is a single scalar value for the entire conversation).The $\Delta\hat{r}_{LI}$ metric evaluates the local predictions for each speaking turn, confirming whether the local predictions are aligned to the local implicit reward. It is normal that the GE-RRD baseline performs well on the first metric, since it is also optimized this way. However, as we observe in the human evaluations and the qualitative visualizations, this GE-RRD baseline ends up being very conservative in its predictions, with little variability in its local predications and often converging to the mean (variance of predicted rewards from GE:RRD is 0.0231 ± 0.004, for GE: RRD+VA is 0.0778 ± 0.006). Hence, it is important to also look at the LI metric where we can observe that for GE:RRD in Table 2 is near 0. Our proposed GELI approach finds a successful balance between both general and local metrics. As we see in the human evaluation in Section 6.2, this GELI balance ends up improving even the widely used LLAMA2 baseline.
@@
6.2 Human Evaluation on Adapted LLM on CANDOR Reece et al. (2023)
We refer the reader to Table 2, where we find that the LLAMA-2 model with GELI outperforms all other approaches in most evaluation metrics and performs comparably with other baselines otherwise. Importantly, if a certain reward function properly adapts the language model via RLHF to become more conversational, it implies that the rewards contains accurate, valuable signals which a reinforcement learning algorithm such as PPO could leverage to improve the policy. For clarity, LLAMA2 + GE, refers to the trained reward function from global explicit reward decomposition only, and LLAMA2 + LI, refers to the trained reward function from local implicit rewards only. Finally, LLAMA2 + GELI refers to our proposed approach, the reward function trained with both global explicit decomposition shaped by local implicit rewards. For dialogue, we find the local implicit rewards (LLAMA2+LI) perform better than that of LLAMA2+GE, where we observe up to a 10% performance boost. However, we find the improvements are often worse than that of the base LLAMA-2 model (3 out of 8 evaluation measures are worse), this leads to the conclusion that the reward signals in GE, and LI separately do not contain enough reward signals to be used as a reward model in a reinforcement learning set-up to adapt the language model to be more conversational. On the otherhand, we find that GELI, by utilizing both GE and LI, gains consistent performance boosts across most conversational evaluation metrics (6 out 8 measures are better, the remaining are comparable), which indicates the combination of both GE and LI contain valuable reward signals for the RL algorithm to utilize.
Overall, compared to base LLAMA-2, we see that there is a significant improvement in the level of emotional connection (+9%), positivity (+18%), understanding of social context (+10%), how interesting the responses are (+14%). It is especially impressive to note that there is a statisical difference in how inclined people wanted to talk to our model over others (+11%), and how much they would want to reuse our chatbot again (+14%). Interestingly, we see statistically signficant results for positivity, which is the most closely related to our primary optimization objective overall-affect, and inclination, reuse, which indicates which chatbot the User would speak to again.
6.3 Generalizability of Adapted LLM on Unseen Dataset: SODA Kim et al. (2023a)
In Fig. 3 we show generalizability of GELI-adapted LLM by running the same experiment and human evaluation from previous Section 6.2 on a new unseen dataset to show generalization. SODA Kim et al. (2023a) is a large social dialogue dataset that was distilled from a social commonsense knowledge graph and generated via GPT 3.5. Human evaluation demonstrates that the dialogue in SODA is more consistent, natural and specific than human-authored datasets. We use the LLAMA2+GELI model trained and CANDOR and evaluate on 100 unseen samples from SODA. We find the GELI performs even better in SODA when compared to CANDOR, performing significantly better results in 7 out of 8 conversational metrics. SODA was generated by ChatGPT, and we find that our proposed approach significantly outperforms ChatGPT by up to 30%. Hence, we can conclude that this approach is generalizable across different datasets and dialogue scenarios.
@@ 7 Conclusion
We introduce GELI, which automatically decomposes a single Global Explicit post-interaction score, incorporating Local Implicit feedback from multimodal behaviors. The reward function trained via GELI is designed to align and improve the conversational capabilities of a language model. GELI performs global alignment of multi-turned interactions by locally rewarding parts of the interaction, shaped by multimodal local implicit feedback. Our proposed approach complements previous alignment approaches, such as RLHF, which requires fine-grained manual reward annotations. We run quantitative and qualitative human studies to evaluate the performance of our GELI approach, with results showing consistent performance boosts across conversational metrics.
@@ Appendix B Human Evaluation Metrics Definitions
The human evaluation scores are conducted via a binary-level classification. For a given question, the annotators can select the models that satisfy the question. For example, for ‘Positivity’, the annotators are given the following question and answer choices:
Which AI responses most likely made User feel positive feelings? (A) (B) (C) (D) (E) (F)
The options A-F refer to models which are randomized in order and anonymized. The annotators can select multiple models if they satisfy the question. Therefore, Table 1 can be interpreted as the percentage of instances out of the samples (300 in our case) where each model satisfied the question.
@@ Appendix F Training Details
F.1 Distribution of GE score (overall-affect):
  • <50: 2.2
  • 50  60: 6.7
  • 60   70: 14.5
  • 70   80: 30.4
  • 80   90: 24.6
  • 90   100: 21.6
Distribution of Emotions Polarity (only Happiness is considered as positive polarity):
  • Anger: 3.9
  • Contempt: 0.08
  • Disgust: 1.98
  • Fear: 2.23
  • Sadness: 8.84
  • Neutral: 35.61
  • Happiness: 40.01
  • Surprise: 7.35